I am looking at timber cladding the gable end of our house, the cladding comes in 5m lengths & is 100mm in height.
My question is:-
How many 5m lengths will I need to order to fill the triangle that is my gable end measuring 7.8m wide at the base & 3.2m high ??
Regards
Steve

Answer by jim_thompson5910(35256)

100 mm = 100/1000 = 0.1 meters


I'm assuming a timber cladding is simply a rectangular board. If so, then each rectangular board has an area of (5 meters)*(0.1 meters) = 5*0.1 = 0.5 square meters


Triangle Area:


A = b*h/2
A = 7.8*3.2/2
A = 12.48


The area of the triangle is 12.48 square meters


-------------------------------------------------------


Each rectangular board is 0.5 square meters
The entire triangle is 12.48 square meters


Let x = number of rectangular boards


(Area of board)*(# of boards) = Area of triangle


0.5x = 12.48


x = 12.48/0.5


x = 24.96


Round up to get 24.96 --> 25. So 25 rectangular boards should fill out an area of 12.48 square meters. Keep in mind that you'll have to make cuts to some of the boards to get them all to fit in the triangular region. So that may require a few more boards if things don't all go smoothly.